# Prompting with Divide-and-Conquer Program Makes Large Language Models Discerning to Hallucination and Deception

### v4

@@ Abstract
Foundation models, such as Large language Models (LLMs), have attracted significant amount of interest due to their large number of applications. Existing works show that appropriate prompt design, such as Chain-of-Thoughts, can unlock LLM’s powerful capacity in diverse areas. However, when handling tasks involving repetitive sub-tasks and/or deceptive contents, such as arithmetic calculation and article-level fake news detection, existing prompting strategies either suffers from insufficient expressive power or intermediate errors triggered by hallucination. To make LLM more discerning to such intermediate errors, we propose to guide LLM with a Divide-and-Conquer program that simultaneously ensures superior expressive power and disentangles task decomposition, sub-task resolution, and resolution assembly process. Theoretic analysis reveals that our strategy can guide LLM to extend the expressive power of fixed-depth Transformer while exploiting smaller context window size compared to existing prompting strategies like CoT. Experiments indicate that our proposed method can achieve better performance than typical prompting strategies in tasks bothered by intermediate errors and deceptive contents, such as large integer multiplication, hallucination detection and misinformation detection.
Machine Learning, ICML
@@ 1 Introduction
Large language models (LLM) based on the Transformer architecture have led to major breakthroughs in natural language processing and other related fields in artificial intelligence(Brown et al., 2020; Radford et al., ; Touvron et al., 2023). State-of-the-art general-purpose language models have demonstrated remarkable advancements in various domains, including question answering, graph learning, reading comprehension, text generation, and machine translation (Chen et al., 2023b; Tan et al., 2023; Hendy et al., 2023; Mao et al., 2023; Zong & Krishnamachari, 2023). These developments paves the way towards general-purpose problem solvers (Bubeck et al., 2023).
@@
However, as pointed out in (Wei et al., 2022), significant challenges arise when scale-up models are applied to tasks involved with long solution paths, such as those requiring mathematical or knowledge reasoning. A series theoretic works attribute this challenge to Parallelism Tradeoff (Merrill & Sabharwal, 2023a), a fundamental limitation of Transformers. Specifically, unlike Recurrent Neural Network whose computational depth is linear to the input sequence length (i.e., the depth is $O(n)$, where $n$ is the input sequence length), Transformer does not contain any recurrent structure. Such design, while achieving superior parallelizability than RNN, makes Transformers suffer from limited expressive power. Merrill & Sabharwal proved that the expressive power of fixed-depth log-precision Transformer, which is very close to the most commonly applied Transformer architecture for LLMs, is bounded by constant-depth logspace-uniform threshold circuits. Thus, they fail to accurately tackle the tasks requiring long solution paths.
@@
Figure 1: An illustrative example of hallucination detection with entangled problem solving (i.e., directly forward all inputs into the LLM) and disentangled problem solving (i.e., divide the problem inputs to parallel sub-tasks and tackle them parallelly). The sentence marked with red back font in the material is the evidence that contradict with the first claim in summary (marked with red font).
To address this challenge, carefully designed prompting strategies have been developed to tackle tasks that requires stronger expressive power (Feng et al., 2023). A series of works focus on prompting the LLM to output the intermediate steps that derive the final answer in an autoregressive manner such as Chain-of-Thoughts (CoT) (Wei et al., 2022; Wang et al., 2022; Yao et al., 2023; Zhou et al., 2022; Chen et al., 2023a). Theoretically, these prompting strategies convert the role of Transformer from the complete problem solver to a sub-problem solver in a dynamic programming or tree searching algorithm (Merrill & Sabharwal, 2023b). In this way, these prompting strategies expand the expressive power of the LLMs and successfully improve the reasoning and searching ability of LLMs (Feng et al., 2023). However, in most of these aforementioned prompting strategies, the processes of sub-problem decomposition, sub-problem resolution, and sub-resolution assembly are intertwined together during the autoregressive token decoding. Such design makes the sub-problem generation process lack control from the human or agent side and susceptible to disruptions of the task resolution process. As a result, when the task requires a large quantity of repetitive tasks (e.g. large integer multiplication and article-level verification), LLM is prone to intermediate errors, such as missing some inputs or generating wrong sub-tasks, leading to problematic answers (Amaro et al., 2023). This phenomenon is especially serious when the task input contains deceptive information or contents that could trigger hallucination (Chen & Shu, 2023; Li et al., 2023a). An illustrative example is presented in the Fig. 1. To alleviate this issue, some program guided prompting strategies such as Least-to-Most (Zhou et al., 2022) (LtM) and Decomposed Prompting (Khot et al., 2022) propose to disentangle sub-task generation and resolution. However, they implement the task decomposer through multi-round conversation (or question-answering) with a LLM to sequentially raise and solve the sub-problems in an alternate manner. When tackling deceptive text like hallucination and fake news, such intertwined process often guides the LLM to sequentially tackle the corpus and thus follow the context’s flow, making LLM prone to deception.
@@
In fact, human brains also suffer from similar hallucination issues (Liu & Sajda, 2023), especially when the tasks are too hard or too complex. For example, when reviewing a long academic paper, some reviewers produce low-quality reviews (Garcia et al., 2021; Tennant & Ross-Hellauer, 2020; Cortes & Lawrence, 2021) containing hallucination-like intermediate errors, such as pointing out some ‘missing baselines’ that have already been sufficiently discussed by authors and requiring the authors to conduct ablation studies for non-existent sub-modules. To avoid such ridiculous mistakes, experienced reviewers usually think slowly (Kahneman, 2011) to follow a Divide-and-Conquer paradigm to handle this task. Specifically, they decompose the paper review as examinations of multiple central opinions. Then they retrieve supporting corpus to verify them respectively. Similar strategies are also applied in decoding to specifically improve LLM’s capacity in math (Lee & Kim, 2023) and automated evaluation (Cui et al., 2023).
Inspired by the aforementioned human experience, in this paper, we explore to guide LLM with Divide-and-Conquer program to unlock the LLM’s ability to handle tasks with repetitive sub-tasks. This strategy breaks down the whole task resolution process into three distinct sub-processes: task decomposition, sub-task resolution, and solution merge. The task decomposition process will prompt the model to separate the whole task to multiple parallelly solvable sub-tasks (i.e., solving one sub-tasks does not require the result of other sub-tasks) and list them explicitly recursively. After that the sub-task resolution process prompts the LLM to output the answer for all sub-tasks. Finally, the solution-merge process assembly the solutions recursively along the decomposition path. The above sub-processes follow a key principle that every upstream sub-process (e.g. sub-task resolution) only forward the final answer to its downstream sub-processes (e.g. solution merge) and does not forward its input and intermediate steps. We denote this principle as disentangled-sub-process principle. An illustrative figure that explain the difference between our method and previous works is provided in Fig. 2
Figure 2: The comparison between our method and the existing methods for prompting. The ellipse marks represent sub-tasks, the right-angled rectangles represent sub-task solutions, and the rounded rectangles represent intermediate steps that entangle sub-task and sub-solutions. The different shades in Tree of Thoughts (subfigure D) indicate the rates of different search directions. In CoT (Chain-of-Thoughts), CoT-SC and ToT, the Large Language Models must simultaneously generating and resolving sub-tasks. Least-to-Most (also Decomposed Prompting) disentangle sub-task generation and resolution. However, as shown in the figure, its sub-task resolution process and resolution assembly process are intertwined as it sequentially attach new sub-tasks onto the previous resolution. Different from all of them, our method totally disentangle the sub-task generation, sub-task resolution, and resolution assembly process.
@@
To validate the expressive power of our proposed prompting strategy, we provide a theoretic analysis to validate that the proposed strategy can expand the expressive power of fixed-depth log-precision Transformer. Meanwhile, the theoretic analysis to computational complexity also reveals that compared to sequential prompting methods like CoT and LtM which can also extend the expressive power, our proposed method generally require smaller context window size for the Transformer (i.e., $O(|t_{b}|)$ vs $O(|t_{p}|\times|t_{b}|)$), alleviating the error propagation in long context (Yang & Ettinger, 2023). To further empirically validate the advantage of our proposed method, we evaluate our method and representative baselines on three tasks that are challenging to existing prompting strategies even on state-of-the-art LLMs: Large Integer Multiplication, Hallucination Detection, Article-level Fact Verification (Cheng & Zhang, 2023; Li et al., 2023a; Wadden et al., 2020; Hu et al., 2023; Wu et al., 2023). These tasks either require very long reasoning paths (e.g. large integer multiplication) or contain deceptive contents (e.g. hallucination detection and fact verification), making existing methods like Chain-of-Thought prompting prone to missing intermediate steps. Meanwhile, these tasks share a common nature that the correctness of the final answer substantially relies on the thoroughness of all steps. Our experimental results show that the proposed method outperforms the baselines on all three tasks.
@@ 2 Related Work
2.1 Limitations of Transformer in Expressive Power
As discussed in previous works (Merrill & Sabharwal, 2023a; Feng et al., 2023), the expressive power of fixed-length log-precision transformers, which are widely applied in modern Pre-trained Large Language Models, is actually much more limited than people’s expects. Merrill & Sabharwal give a theoretic proof that the expressive power of fixed-length log-precision transformers is upper-bounded with ${\sf TC^{0}}$. Feng et al. further extend their analysis to explain that a lot of common problems, such as calculating arithmetic expressions, exceeds the expressive power of fixed-length log-precision transformers. Such results explains why the powerful Large Larnguage Models may make some ridiculous mistakes and why CoT can improve the performance.
2.2 Prompting Strategies for Large Language Models
@@
Input-Output (IO) Prompting is the standard prompting strategy that attach input $x$ with instructions and/or few-shot in-context-learning examples to aqcuaire a prompt, denoted as ${\sf prompt}(x)$ (Yao et al., 2023). The LLM takes ${\sf prompt}(x)$ as input and predict result, i.e. $y\sim p_{\theta}(y|{\sf prompt}(x))$. The main drawback of IO prompting is its poor expressive power. Existing works have shown the expressive power limitation of IO prompting. (Merrill & Sabharwal, 2023a) proves that log-precision transformer based LLM equipped with prompt of polynomial length is a subset of ${\sf TC^{0}}$ class.
@@
Exploration-of-Thought (EoT) Prompting includes a series of CoT’s variants, such as Self-consistency with CoT (CoT-SC) prompting (Wang et al., 2022) and Tree-of-Thoughts (ToT) prompting (Yao et al., 2023), which aim at addressing the limitation of CoT in exploration. Their common central idea is to generate multiple chains of thought through sampling or proposing prompting and then ensemble them to acquire a final prediction.
@@
Program-guided Prompting aims at controlling the LLM’s generation process with symbolic programs or pre-defined procedure (Zhu et al., 2022; Jung et al., 2022; Zhou et al., 2022; Khot et al., 2022; Creswell & Shanahan, 2022; Gao et al., 2023). Among them, the Least-to-Most (LtM) Prompting (Zhou et al., 2022) and Decomposed Prompting (Khot et al., 2022) are close to this work. They are the earliest attempts that explicitly prompt the LLM to decompose the task as a series of sub-tasks and sequentially tackle them. LtM prompt a LLM to iteratively raise sub-tasks and sequentially solve them to acquire the final resolution. Decomposed Prompting can regarded as a upgraded version of LtM. It introduces special notations into the prompt to represent program states and thus can call itself (i.e., recursion) or other modules (i.e., hierarchical decomposition), endowing it stronger expressive power. Such design increased the compositional generalization ability of LLMs in different areas, such as symbolic manipulation and multi-hop QA (Khot et al., 2022).
@@
The aforementioned CoT and EoT families incorporate LLM with stronger expressive power than IO prompting. However, a critical issue of them is that, they could miss or ignore some important intermediate steps or contents (Liu et al., 2023). This problem is even worse when we are handling tasks involved with long input (e.g. long documents and large numbers). Typical examples include large number arithmetic calculation and fact verification in long documents. Compared to them, Least-to-Most prompting and Decomposed Prompting introduce explicit task decomposition to enumerate sub-tasks. However, their task decomposers are based on multi-round conversation or question-answering, which often guides the LLM to sequentially tackle the sub-tasks. When tackling contents containing deceptive contents, e.g. hallucination and fake news detection, such design has two issues. First, it will navigate the LLM through the deceptive content’s flow and thus make LLM prone to deception. Second, as pointed by previous works (Yang & Ettinger, 2023; Dziri et al., 2024), when the round of conversation increases, the context window size for auto-regressive decoding also increases, making the hallucination/deception propagate and accumulate in the contexts, leading to decreased performance and accuracy.
3 Proposed Method
@@
To avoid the task decomposition and task resolution from interweaving and interrupting each other, we propose to guide LLM with Divide-and-Conquer (DaC) program that consists of three distinct stages: task decomposition stage, sub-task resolution stage, solution merge stage. In task decomposition stage, the LLM is prompted to explicitly decompose the task as a series of parallel homogeneous sub-tasks with smaller problem sizes (e.g. divide a long paragraph to sentences). Such parallel design ensures LLMs can tackle sub-tasks without relying on a specific resolution order. As a result, the correctness of each sub-task’s resolution does not rely on other sub-tasks, making LLMs less prone to intermediate errors and deceptions, and thus the issues of multi-round conversation in LtM and Decomposed Prompting can be avoided. After that, in sub-task resolution stage, the LLM is prompted to provide the solutions for every sub-task. Finally, in the solution merge stage, the LLM is prompted to assembly the solutions of subtasks and acquire the final answer. In this process, all three stages are isolated to avoid interruption. They are all guided by a program rather than a LLM to avoid hallucination or the deception from input context. To tackle tasks of different sizes, we propose two variants: Single-Level DaC Solver and Multi-Level DaC Solver.
@@
Single-level Divide-and-Conquer Solver decomposes the task in one call to the LLM, which expands the original task as a tree of one level. The algorithm is presented in the Alg. 1. The advantage of this variant is its simplicity and efficiency. However, when the original input is too long, single-level Divide-and-Conquer Solver may acquire sub-tasks with large problem sizes that will still trigger intermediate errors. In such a case, following (Khot et al., 2022; Lee & Kim, 2023), we can recursively expand the task as a multi-level tree. More specifically, we repeat the aforementioned steps to further divide the sub-tasks hierarchically until they are easy enough to be handled by the LLM. This can be done through a recursion program as presented the Alg. 2.
@@
4 Theoretic Analysis on the Proposed Method
In this section, we provide a theoretic analysis to indicate that the proposed Divide-and-Conquer program-guided problem solving strategy help expand the expressive capacity of Transformers. Specifically, we exploit 2-color Binary Subtree Isomorphism (2-BSI) problem for the proof. 2-BSI problem is a ${\sf NC^{1}}$-complete problem. Under a widely-accepted assumption in parallel computing, we theoretically prove two facts: (1) Log-Precision Transformers with standard IO-prompting strategy can not solve 2-BSI problem, and (2) there exists a Log-Precision Transformers with Divide-and-Conquer program-guided problem solving strategy that can solve 2-BSI problem. In this way, we prove that Divide-and-Conquer strategy expand the expressive power of Large Language Models.
4.1 Fixed-Depth Transformer can not solve 2-color Binary Subtree Isomorphism (2-BSI)
In this section, we will present a problem that exceeds the expressive power of fixed-length log-precision transformers: 2-color Binary Subtree Isomorphism (also known as tree matching), an important algorithmic problem which many tasks such as semantic similarity analysis, text matching and structured text database querying can be reduced to (Kilpeläinen & Mannila, 1992; Marsi & Krahmer, 2010; He et al., 2018). More specifically, we first gives a theoretical proof that 2-color Binary Subtree Isomorphism is not solvable for fixed-length log-precision transformers.

### v6

# An Examination on the Effectiveness of Divide-and-Conquer Prompting in Large Language Models

@@ Abstract
Foundation models, such as Large language Models (LLMs), have attracted significant amount of interest due to their large number of applications. However, when handling tasks involving repetitive sub-tasks and/or deceptive contents, such as arithmetic calculation and article-level fake news detection, simple instructional prompts suffer from inaccurate responses. Existing works show that more complicated prompting strategies, such as Chain-of-Thoughts and Least-to-Most, can unlock LLM’s powerful capacity in diverse areas. Recent researches reveal that simple divide-and-conquer prompting strategy, i.e. simply dividing the input sequence to multiple sub-inputs, can also substantially improve LLM’s performance in some specific tasks such as misinformation detection. In this paper, we aim at examining the utility of divide-and-conquer prompting strategy and answer on which kind of tasks this strategy gets advantages. Specifically, we provide a theoretic analysis to divide-and-conquer prompting strategy and help us identify the specific tasks where DaC prompting can bring performance boost with theoretic guarantee. We then present two cases (large integer arithmetic and fact verification) where experimental results aligns with our theoretic analysis.
An Examination on the Effectiveness of Divide-and-Conquer Prompting in Large Language Models
Yizhou Zhang1, Lun Du2, Defu Cao1, Qiang Fu3, Yan Liu1 1 University of Southern California; 2 Coupang, Inc; 3 Microsoft Research, Asia {zhangyiz, defucao, yanliu.cs}@usc.edu dulun2834@126.com, qifu@microsoft.com
@@ 1 Introduction
Large language models (LLM) based on the Transformer architecture have led to major breakthroughs in natural language processing and other related fields in artificial intelligence(Brown et al., 2020; Radford et al., ; Touvron et al., 2023). State-of-the-art general-purpose language models have demonstrated remarkable advancements in various domains, including question answering, graph learning, reading comprehension, text generation, and machine translation (Chen et al., 2023b; Tan et al., 2023; Hendy et al., 2023; Mao et al., 2023; Zong and Krishnamachari, 2023). These developments paves the way towards general-purpose problem solvers (Bubeck et al., 2023).
@@
However, as pointed out in (Wei et al., 2022), significant challenges arise when scale-up models are applied to tasks involved with long solution paths, such as those requiring mathematical or knowledge reasoning. A series theoretic works attribute this challenge to Parallelism Tradeoff (Merrill and Sabharwal, 2023b), a fundamental limitation of Transformers. Specifically, unlike Recurrent Neural Network whose computational depth is linear to the input sequence length (i.e., the depth is $O(n)$, where $n$ is the input sequence length), Transformer does not contain any recurrent structure. Such design, while achieving superior parallelizability than RNN, makes Transformers suffer from limited expressive power. Merrill and Sabharwal proved that the expressive power of fixed-depth log-precision Transformer, which is very close to the most commonly applied Transformer architecture for LLMs, is bounded by constant-depth logspace-uniform threshold circuits. Thus, they fail to accurately tackle the tasks requiring long solution paths.
@@
Figure 1: An illustrative example of hallucination detection with entangled problem solving (i.e., directly forward all inputs into the LLM) and divide-and-conquer problem solving (i.e., divide the problem inputs to parallel sub-tasks and tackle them parallelly). The sentence marked with red back font in the material is the evidence that contradict with the first claim in summary (marked with red font).
To address this challenge, carefully designed prompting strategies have been developed to tackle tasks that requires stronger expressive power (Feng et al., 2023). A series of works focus on prompting the LLM with instructions or context samples to output the intermediate steps that derive the final answer in an autoregressive manner, such as Chain-of-Thoughts (CoT) (Wei et al., 2022; Wang et al., 2022; Zhou et al., 2022; Chen et al., 2023a). Some works further apply programs to guide LLM to strictly follow designated reasoning steps (Yao et al., 2023). Theoretically, these prompting strategies convert the role of Transformer from the complete problem solver to a sub-problem solver in a dynamic programming or tree searching algorithm (Merrill and Sabharwal, 2023a). In this way, these prompting strategies expand the expressive power of the LLMs and successfully improve the reasoning and searching of LLMs (Feng et al., 2023).
@@
In contrast to such methods that apply instruction, context sample or program to decompose the whole reasoning process to multiple intermediate steps, in some tasks, researchers report that LLM’s performance can also be boosted by simply dividing the input sequences to multiple sub-inputs and then merge the responses from LLMs on all sub-inputs, as shown in Fig. 1. For example, Cui et al. propose that in automated evaluation, LLM’s performance can be further boosted by first dividing the input text to sentences and then evaluating them one by one. Intuitively, this paradigm benefits the tasks in a way similar to human brains, especially when the tasks are too hard or too complex. For example, when reviewing a long academic paper, some reviewers produce low-quality reviews (Garcia et al., 2021; Tennant and Ross-Hellauer, 2020; Cortes and Lawrence, 2021) containing hallucination-like intermediate errors, such as pointing out some ‘missing baselines’ that have already been sufficiently discussed by authors. To avoid such mistakes, experienced reviewers usually think slowly (Kahneman, 2011) to follow a Divide-and-Conquer paradigm to handle this task. Specifically, they decompose the paper review as examinations of multiple central opinions and then retrieve corpus to verify them respectively.
However, unlike Chain-of-Thoughts whose advance in expressive power is supported by theoretic analysis (Feng et al., 2023), the performance boost from Divide-and-Conquer paradigm is lack of rigorous theoretic support. As a result, we are not aware of the conditions under which the Divide-and-Conquer paradigm can acquire more accurate answers. To tackle this challenge, in this paper, we aim at understanding the utility of DaC prompting. More specifically, we attempt to answer the following two research questions:
  1. 1.
    RQ1: Compared to straightforward instructional prompting, does DaC have theoretically guaranteed advantages similar as CoT and its variants?
  2. 2.
    RQ2: Compared CoT and its variants, what utility and limitations does DaC have?
@@
To answer these questions, we first provide a theoretic paradigm that can help us analyze how divide-and-conquer strategy expand the expressive power of fixed-depth log-precision Transformer on a given task. In this way, we provide a framework that can provide theoretic guarantee to DaC paradigm in various tasks. In this way, we present some conditions under which DaC have advantages compared to other prompting strategies. We then empirically evaluate DaC prompting and representative baselines on tasks that satisfy the proposed conditions and are challenging to existing prompting strategies even on state-of-the-art LLMs: Large Integer Multiplication, Hallucination Detection, Article-level Fact Verification (Cheng and Zhang, 2023; Li et al., 2023a; Wadden et al., 2020; Hu et al., 2023; Wu et al., 2023). These tasks either require very long reasoning paths (e.g. large integer multiplication) or contain deceptive contents (e.g. hallucination detection and fact verification), making existing methods like Chain-of-Thought prompting prone to intermediate errors. Our experimental results show that the proposed method outperforms the baselines on all three tasks, which supports our theoretic analysis.
Figure 2: The comparison between DaC and the existing methods for prompting. The ellipse marks represent sub-tasks, the right-angled rectangles represent sub-task solutions, and the rounded rectangles represent intermediate steps that entangle sub-task and sub-solutions. The different shades in Tree of Thoughts (subfigure D) indicate the rates of different search directions. In CoT (Chain-of-Thoughts), CoT-SC and ToT, the Large Language Models must simultaneously generating and resolving sub-tasks. Least-to-Most (also Decomposed Prompting) disentangle sub-task generation and resolution. However, its sub-task resolution and resolution assembly process are intertwined as it sequentially attach new sub-tasks onto the previous resolution. Different from them, DaC totally disentangle the sub-task generation, sub-task resolution, and resolution assembly process.
@@ 2 Related Work
2.1 Expressive Power of Transformer
As discussed in previous works (Merrill and Sabharwal, 2023b; Feng et al., 2023), the expressive power of fixed-length log-precision transformers, which are widely applied in modern Pre-trained Large Language Models, is actually much more limited than people’s expects. Merrill and Sabharwal give a theoretic proof that the expressive power of fixed-length log-precision transformers is upper-bounded with ${\sf TC^{0}}$. Feng et al. further extend their analysis to explain that a lot of common problems exceed the expressive power of fixed-length log-precision transformers. Such results explains why the powerful LLM may make some ridiculous mistakes and how CoT improve the performance.
2.2 Prompting Strategies of LLM
@@
Input-Output (IO) Prompting is the standard prompting strategy that attach input $x$ with instructions and/or few-shot in-context-learning examples to aqcuaire a prompt, denoted as ${\sf prompt}(x)$ (Yao et al., 2023). The LLM takes ${\sf prompt}(x)$ as input and predict result, i.e. $y\sim p_{\theta}(y|{\sf prompt}(x))$.
@@
Exploration-of-Thought (EoT) Prompting and Program-guided Prompting are two variants of CoT. EoT includes a series of CoT’s variants, such as Self-consistency with CoT (CoT-SC) prompting (Wang et al., 2022) and Tree-of-Thoughts (ToT) prompting (Yao et al., 2023), which aim at addressing the limitation of CoT in exploration. Their common central idea is to generate multiple chains of thought through sampling or proposing prompting and then ensemble them to acquire a final prediction. Program-guided Prompting aims at controlling the LLM’s generation process with symbolic programs or pre-defined procedure (Zhu et al., 2022; Jung et al., 2022; Zhou et al., 2022; Khot et al., 2022; Creswell and Shanahan, 2022; Gao et al., 2023). Among them, the Least-to-Most (LtM) Prompting (Zhou et al., 2022) and Decomposed Prompting (Khot et al., 2022) are close to this work. They are the earliest attempts that explicitly prompt the LLM to decompose the task as a series of sub-tasks and sequentially tackle them. LtM prompt a LLM to iteratively raise sub-tasks and sequentially solve them to acquire the final resolution. Decomposed Prompting can regarded as a upgraded version of LtM. It introduces special notations into the prompt to represent program states and thus can call itself (i.e., recursion) or other modules (i.e., hierarchical decomposition), endowing it stronger expressive power. Such design increased the compositional generalization ability of LLMs in different areas, such as symbolic manipulation and multi-hop QA (Khot et al., 2022).
@@
The aforementioned CoT and EoT families incorporate LLM with stronger expressive power than IO prompting. However, a critical issue of them is that, they could miss or ignore some important intermediate steps or contents (Liu et al., 2023). This problem is even worse when we are handling tasks involved with long input (e.g. long documents and large numbers). Typical examples include large number arithmetic calculation and fact verification in long documents. Compared to them, Least-to-Most prompting and Decomposed Prompting introduce explicit task decomposition to enumerate sub-tasks. However, their task decomposers are based on multi-round conversation or question-answering, which navigate the LLM through the deceptive content’s flow sequentially, and propagate the hallucination/deception in the contexts (Dziri et al., 2024; Yang and Ettinger, 2023), leading to decreased performance.
3 Preliminary of Divide-and-Conquer Prompting
@@
In this section, we summarize and formalize Divide-and-Conquer prompting strategy. Divide-and-Conquer prompting strategy consists of three distinct stages: task decomposition stage, sub-task resolution stage, solution merge stage. In task decomposition stage, the LLM is prompted to explicitly decompose the task as a series of parallel homogeneous sub-tasks with smaller problem sizes (e.g. divide a long paragraph to sentences). Such design avoids the multi-round conversation or question-answering in LtM and Decomposed Prompting, making the model less prone to deception. After that, in sub-task resolution stage, the LLM is prompted to provide the solutions for every sub-task. Finally, in the solution merge stage, the LLM is prompted to assembly the solutions of subtasks and acquire the final answer. To tackle tasks of different sizes, Divide-and-Conquer prompting strategy can be divided to two variants: Single-Level DaC Solver and Multi-Level DaC Solver.
@@
Single-level Divide-and-Conquer Solver decomposes the task in one call to the LLM, which expands the original task as a tree of one level. The algorithm is presented in the Alg. 1. The advantage of this variant is its simplicity and efficiency. However, when the original input is too long, single-level Divide-and-Conquer Solver may acquire sub-tasks with large problem sizes that will still trigger intermediate errors. In such a case, following (Khot et al., 2022), we can recursively expand the task as a multi-level tree. More specifically, we repeat the aforementioned steps to further divide the sub-tasks hierarchically until they are easy enough to be handled by the LLM. This can be done through a recursion program as presented in Alg. 2. More discussions on the proposed method’s appliance scope, including its comparison with other prompting strategies and limitations, can be found in A.1
@@
4 Main Theoretic Results
In this section, we provide theoretic analysis to the utility and limitations of the Divide-and-Conquer prompting. In the first subsection, we provide a comparison of IO prompting (common fixed-length instructional prompting) and DaC prompting in expressive power perspective. This part answers the first research question: the expressive power of IO prompting is a subset of DaC prompting. In the second subsection, we provide a comparison between Chain-of-Thoughts and DaC prompting in expressive power. Our comparison suggests that, although the expressive power of DaC prompting is a subset of Chain-of-Thoughts, for tasks satisfying specific conditions, DaC prompting can solve the problem with lower average context window length when decoding the tokens. Such property is empirically proved to be helpful for reducing the intermediate error and thus boost the performance.
4.1 Divide-and-Conquer vs. IO Prompting
We show that the expressive power of Divide-and-Conquer is stronger than IO Prompting:
Theorem 4.1.
We denote the set of problems that a fixed-precision transformer with fixed-length IO prompting can tackle as $S(IO)$. Similarly, we denote the set of problems that a fixed-precision transformer with DaC prompting can tackle as $S(DaC)$. Then we have the following results:
$S(IO)\subset{\sf TC^{0}}\subseteq{\sf NC^{1}}\subseteq S(DaC)$ (1)
Proof Sketch: The conclusion that $S(IO)\subset{\sf TC^{0}}$ is a corollary of the main results in (Chiang et al., 2023). In this paper, we mainly focus on proving ${\sf NC^{1}}\subseteq S(DaC)$. Specifically, we exploit 2-color Binary Subtree Isomorphism (2-BSI) problem for the proof. In (Jenner et al., 2003), 2-BSI problem is proved to be an ${\sf NC^{1}}$-complete problem. Its definition is: